A study looking at symptoms and care-seeking behavior in African-American patients with chest pain, measured the systolic blood pressure of 157 African-American men seen in an emergency department. The mean systolic blood pressure was 146 mm Hg with a standard deviation of 27. The researchers wish to know if on the basis of these data, they can conclude that the mean systolic blood pressure is greater than 140. Use α=.05.
Answer #1

H0: <= 140

Ha: > 140

Test statistics

t = - / (S / sqrt(n) )

= 146 - 140 / (27 / sqrt(157) )

= 2.78

This is test statistics value.

df = n -1 = 157 - 1 = 156

t critical value at 0.05 level with 156 df = 1.655

Since test statistics value falls in rejection region , that is greater than 1.655, we have sufficient

evidence to reject H0.

We conclude at 0.05 level that we have enough evidence to support the claim.